# Controlled Training Data Generation with Diffusion Models

### v1

@@ 4.2 Depth Estimation
Performance of GAP against amount of generated data. Fig. 5-right shows the performance of our method on the defocus blur corruption over the amount of extra generated data (See Appendix Fig. 18 for results on other corruptions). Guided Prompts or Guided Adversarial Prompts results in a large improvement in performance, compared to Adversarial Prompts or the baseline with only 10 extra data points. This suggests that the guidance loss successfully steered the generations toward producing training data relevant to the distribution shift. This experiment was performed with image guidance. In Appendix Sec. 8.4.2, we compare the qualitative and quantitative differences from image and text guidance. Text guided prompts tends to generate corruptions that are more realistic (see Appendix Fig. 18 for comparisons). However, image guidance tends to perform better quantitatively across the different target distributions.
Unlike classification, the performance of Guided Adversarial Prompts for depth is not consistent across all distribution shifts. The diffusion model was not able to generate certain shifts e.g., noise corruptions, as the text descriptions and unlabelled images was too ambiguous e.g., ‘noise’ or having common attributes other than the corruption. We leave further analysis to future work.
@@
Running multiple iterations of adversarial optimization vs a single iteration. We define an iteration as one round of adversarial optimization, i.e. optimizing Eq. 1 or Eq. 3, generation and fine-tuning. Given that all of the above results were obtained with a single iteration, we aim to see if there are benefits in performing multiple iterations. We perform a total of 8 iterations and we compare this to performing a single iteration. The experimental settings for 8 iterations and a single iteration are similar, e.g., in total, over the 8 iterations, we optimize for the same number of Adversarial Prompts, and fine-tune on the same datapoints, etc. Fig. 5(b) shows that the first iteration of the 8 iterations setting resulted in the largest improvement in performance, eventually converging to the performance of the single iteration approach. Thus, we chose to perform a single iteration for the results in Fig. 5(a). See the Appendix Sec. 8.4.1 for further implementation details and results.
@@ 5 Conclusion and Limitations
In this work, we aim to generate training data useful for training a supervised model by steering a text-to-image generative model. We introduced two feedback mechanisms to find prompts that are informed by both the given model and the target distribution. Evaluations on a diverse set of tasks and distribution shifts show the effectiveness of the proposed closed-loop approach in comparison to open-loop ones. Below we briefly discuss some of the limitations:
  • Label shift: In this work, we focus on generating novel images. However, some distribution shifts can also change the label distribution, e.g., for depth estimation, changing from indoor to outdoor scenes would result in a shift in depth maps. One possible approach could be learning a generative model over the label space [43] to control the generation in both the label and image space.
  • Computational cost: Estimating the gradient of the loss in Eq. 3 requires backpropagation through the denoising process of the diffusion model, which can be computationally demanding. Using approaches that reduce the number of denoising steps [71, 48] may be able to reduce this computational cost.
@@ 6 Outline
  • Secs. 7 and 8 describe additional implementation details for our classification and depth estimation experiments, respectively.
  • Sec. 7.6 describes an image guidance mechanism using Textual Inversion [22] and compares it with the CLIP guidance mechanism, on Waterbirds.
  • Sec. 8.2 provide additional results for “standard” augmentation baselines for the depth estimation experiments.
  • Secs. 7.4.3, 7.5.4 and 8.3 provide qualitative generations from all the Adversarial Prompts and Guided Adversarial Prompts used in the Waterbirds, iWildCam, and depth estimation experiments.
    • Additionally, for depth estimation, we provide a qualitative comparison of Adversarial Prompts generations optimized on different models (UNet [63], DPT [58]). For iWildCam, we also provide additional results using a ViT-B-16 [17] instead of a ResNet50 [26].
  • Sec. 8.4 provides additional analysis on the depth estimation experiments:
@@
7 Classification
7.1 Training data generation
@@
As mentioned in main paper Sec. 3.1, for semantic classification tasks, we utilize the foreground object masks and use an in-painting technique proposed in [47] that preserves the masked region throughout the denoising process. In this section, we briefly describe this procedure and refer the reader to the original work for more details.
@@
However, because we are using Stable Diffusion [61], the denoising process is done in latent space (using an encoder $\mathcal{E}$), not pixel space. This means that to apply inpainting, we must resize the mask $m$ to the latent space dimensions, and apply the above-described procedure in the latent space: $\tilde{z}_{t}\leftarrow m_{z}\odot z^{\mathrm{orig}}_{t}+(1-m_{z})\odot\tilde{z}_{t}$, where $z^{\mathrm{orig}}_{0}=\mathcal{E}(x^{\mathrm{orig}})$ and $z^{\mathrm{orig}}_{t}$ is its corresponding noised version. While this procedure usually performs well in preserving the original region of interest, we also paste the original masked region in the pixel space to obtain the final sample $\tilde{x}=m\odot x^{\mathrm{orig}}+(1-m)\odot\tilde{x}_{0}$.
SDEdit [50].
In addition to inpainting, depending on the setting, we also use SDEdit [50], a mechanism available to all diffusion models that allows to use an initial image to condition the generation of new images to be closer to the initial image. The mechanism is parametrized by the SDEdit strength $s$, which indicates the extent by which the model can deviate from the original image.
@@
7.2 ALIA
Here, we give more details on the ALIA [18] (Automated Language-guided Image Augmentation) baseline method, which aims at generating images targeting a particular test distribution similar to our guidance mechanism (main paper Sec. 3.3).
Given exemplar images from the test distribution, ALIA first captions each image using the BLIP [45] captioning model. Then, it uses the GPT-4 [56] LLM to summarize these captions into a list of domains asking it to produce descriptions that are agnostic to the class information. [18] then use these prompts to generate additional training data. In order to preserve the original class information in their generations, they use SDEdit [50] or Instruct Pix2Pix [7]. We refer the original paper for further implementation details. Below, we summarize resulting prompts we use for comparison in our results.
For Waterbirds [66], we found that removing the prefix “a photo of a {class name} from the original prompts when using the inpainting technique (Sec. 7.1) to work slightly better for both the ALIA baseline and our CLIP text guidance (main paper Eq. 2). We, therefore, use the following prompts:
@@ 7.2 ALIA
For iWildCam [3], we keep the original prompts intact:
@@
7.3 General implementation details
@@
We report our generation parameters in Tab. 3. We use the DDIM [71] scheduler. We generate 384x384 resolution images. Those parameters were chosen based on visual inspection, ease of optimization and downstream performance (validation accuracy). Training data. After generation, ALIA’s method consists of an additional filtering step to remove “bad” generations. This step relies on using a pretrained model to measure confidence on the generated images. However, given our method creates images that are adversarial to an iWildCam pretrained model, the filtering part of ALIA’s pipeline is not usable on our data. Thus, to keep things comparable, we decided not to apply filtering both our method generated data and ALIA’s generated data. However, it must be noted that [18] only reports a 2% absolute accuracy drop between fitlering and no filtering on iWildCam (1.4% on Waterbirds), thus we do not expect a big difference in performance with ALIA’s reported results and our results.
@@
We report our training parameters in Tab. 4. We use ALIA’s codebase to finetune our models, which ensures fair comparison to the ALIA baselines. For everything except the generated data, the settings are the same as in ALIA. For both datasets, the starting model is a ResNet50 [25] model, pretrained on ImageNet [13].
@@
7.4 Waterbirds
7.4.1 Dataset details
Fig. 7 demonstrates the shift between train and test distributions in the Waterbirds dataset [65]. We follow the setting suggested in [18] and use 1139 images as $\mathcal{D}_{\mathrm{tr}}$, where waterbirds appear only on water background and landbirds on land background. We add additional 839 examples either from the original dataset, where waterbirds appear only on land background and landbirds on water background (”Real OOD data”), or generated by Stable Diffusion with prompts obtained by one of the methods. For the data-efficiency plots (e.g., Fig. 3) we reduced the number of added examples by a factor of {1/2, 1/4, 1/8, 1/16}.
Since the original Waterbirds dataset does not provide masks for the exact generated images, we used the SAM [40] segmentation model to obtain bird segmentation masks for training images. We use these masks to condition the generative model on the class by using inpainting as described in Sec. 7.1.
7.4.2 Implementation Details
@@
For adversarial feedback, we use the model trained only using the original training data $\mathcal{D}_{\mathrm{train}}$ with complete spurious correlation. It is taken from ALIA checkpoints [^3] https://api.wandb.ai/files/clipinvariance/ALIA-Waterbirds/y6zc932x/checkpoint/ckpt-Waterbirds-none-filtered-resnet50-1-0.001-0.0001/best.pth. As the task is the binary classification, we use the cross-entropy loss for the opposite class as the adversarial loss: $\mathcal{L}_{\mathrm{adv}}(\tilde{x},y)=\mathcal{L}_{\mathrm{x-ent}}(f(\tilde{x}),1-y)$, assuming $y\in\{0,1\}$. This is equivalent to the negative cross-entropy loss referred in the text. We find four prompts per each class, i.e., eight prompts in total. Each prompt is composed of five new learnable tokens. We perform adversarial optimization for 1000 steps with learning rate 1e-3 using Adam [38]. We use five denoising steps during adversarial optimization and generate images for training with 15 steps. We do not use SDEdit for Waterbirds. See Tab. 3 for summary.
@@
For Waterbirds, we use CLIP text guidance by encoding each of ALIA’s summarized prompts (see Sec. 7.2) with the CLIP text encoder as described in main paper Sec. 3.3. In addition, we renormalize the averaged target text embedding to have the norm equal to the mean norm of the original prompts, and use the resulting vector as the target $e_{\mathrm{t}}$. We use $l_{2}$ guidance loss: $\mathcal{L}_{\mathrm{t}}(E_{\mathrm{t}}(c_{w}),e_{\mathrm{t}})=\|E_{\mathrm{t}}(c_{w})-e_{\mathrm{t}}\|_{2}^{2}$. We use $\lambda_{\mathrm{t}}=20$ and $\lambda_{\mathrm{i}}=0$ (i.e., no image guidance).
7.4.3 Additional Qualitative Results.
@@
7.5 iWildCam
7.5.1 Dataset details.
The original iWildCam [3] dataset is subsampled to create a 7-way classification task (background, cattle, elephant, impala, zebra, giraffe, dik-dik). The training set has 6,000 images with some classes having as few as 50 images per example. There are 2 test locations that are not in the training or validation set. Additionally, given $h$, the hour at which an image was taken, we define an image to be during “daytime” if $9\leq h\leq 17$, and “nighttime” if $h\leq 5\lor h\geq 20$. As said in main paper Sec. 4.1, for image CLIP guidance, we separate the target test locations into four groups (location={1,2}, time={daytime, nighttime}). We provide visualisation of the test locations (at day & night) in Fig. 10. For more details on the iWildCam subset construction, we refer to [18] Section 8.3. For inpainting, the object masks are obtained from MegaDetector [2].
@@
7.5.2 Alignment collapse solution for iWildCam.
@@ 7.5.3 Implementation details
We use the “entropy” loss, described previously. For adversarial prompts, we train for 2000 steps. For guided adversarial prompts, we use CLIP guidance coefficient with $\lambda_{i}=10$ and and $\lambda_{\mathrm{t}}=0$ (i.e., no text guidance). We train for a total of 10000 gradient steps. However, we don’t optimize the adversarial loss for the first 2000 steps to allow the prompt to first converge to the target distribution region.
@@
The generation settings are the same as the ones used during adversarial optimization. For each target domain-guided adversarial prompt, (i.e. location & time of the day), the source images (used to condition the generation with an object mask and through SDEdit [50]) are only images that match the time of the day of the target domain used during generation. Furthermore, for each prompt, we only generate one image per source image.
@@
Our iWildCam pretrained model is taken from ALIA checkpoints [^5] https://api.wandb.ai/files/clipinvariance/ALIA-iWildCamMini/brr7b3ks/checkpoint/ckpt-iWildCamMini-randaug-filtered-resnet50-0-0.001-0.0001/best.pth. ALIA trains the model from “scratch” (i.e. the model has never seen iWildCam data), for 100 epochs, on the combination of real + generated data. For our method, given we optimize the prompts based on a finetuned model feedback, it may not make as much sense to train the model from ”scratch”. Thus, we also introduce the variant where the iWildCam pretrained model is finetuned on the combination of real + generated data for 20 epochs, where finetuning means that every layer, except the last, is frozen.
@@
7.5.4 Additional Qualitative Results.
@@
7.5.5 Using ViT model.
In Fig. 14, we repeat the iWildCam experiment from the main paper (Fig. 4) with a ViT-B-16 [16] model. Additionally, we provide qualitative results for generations from adversarial prompts optimized with a ViT-B-16 model in Fig. 15.
@@
7.6 Image Guidance using Textual Inversion
In addition to the CLIP image guidance introduced in main paper Sec. 3.3, we also explore using Textual Inversion (TI) [22] as an image guidance mechanism. Similar to the CLIP guidance, we use a few images $\{x_{j}\}$ from the target distribution. Now, instead of the similarity in a CLIP embedding space, we use the denoising loss between a generated image and one of the target images as in [22] (see Eq. (2)):

### v2

@@ 4.2 Depth Estimation
Quantitative results. Tab. 1 shows the results from finetuning our method and the baseline on the different generated datasets. Adversarial Prompts improves the performance of the model under different distribution shifts for both the UNet and DPT models. Fig. 5 (Right) shows the results with Guided Adversarial Prompts for the defocus blur corruption against the amount of extra generated data used for finetuning. Guided Prompts and Guided Adversarial Prompts result in a large performance improvement, compared to Adversarial Prompts or the baseline with only 10 extra data points. This suggests that the guidance loss successfully steered the generations toward producing training data relevant to the distribution shift. See Fig. 18 for results on other corruptions. As the dataset size increases, GP shows signs of saturation, whereas GAP further improves the model, showing the benefit of the adversarial feedback mechanism.
4.3 Additional Analysis
Table 2: Transferability of adversarial prompts. We show the results of different combinations of the finetuning model and model used to generate additional data via model-informed feedback. We report top-1 accuracy for iWildCam and $l_{1}$ loss for depth.
iWildCam, Acc. ($\uparrow$) Depth, $l_{2}$ ()($\downarrow$)( ↓ )
[…]
ResNet 83.97 72.61 UNet 4.73 3.39
ViT 83.87 77.21 DPT 4.88 3.46
Agnostic
Prompts
72.94 73.07
Agnostic
Prompts
5.27 3.66
Are adversarial prompts found for one model specialized for that model? We study 1) whether models benefit the most from data generated using their own feedback and 2) whether they benefit at all from using data generated for another model (compared to Agnostic Prompts.) Tab. 2 shows that in most cases, a model performs best when trained with data generated using its own feedback, which shows the usefulness of the model-informed feedback mechanism. At the same time, we find that in most cases, prompts found for another model still outperform Agnostic Prompts, suggesting that adversarial prompts result in useful transfer between models.
Figure 6: Comparing the performance from running multiple iterations versus a single iteration of adversarial optimization, generation, and fine-tuning. The plot shows the $\ell_{1}$ loss (×$\times 100$) of the U-Net model against the number of iterations on the depth prediction task. The loss is computed on the Taskonomy dataset under common corruptions, averaged over all corruptions and severity levels.
@@ 4.3 Additional Analysis
Does multiple iterations of adversarial optimization further improve performance? We define an iteration as one round of adversarial optimization, i.e. optimizing Eq. 1 or Eq. 3, generation and fine-tuning. Given that all of the above results were obtained with a single iteration, we aim to see if there are benefits in performing multiple iterations. We perform a total of 8 iterations keeping the total number of generated images, prompts, etc., the same as for a single iteration for comparison. Fig. 6 shows that the first iteration of the multi-iterations setting results in the largest performance improvement and eventually converges to the performance of the single iteration approach. Thus, we chose to perform a single iteration for our main experiments. See the Sec. C.4.1 for additional implementation details and results.
@@ 5 Conclusion and Limitations
In this work, we aim to generate training data useful for training a supervised model by steering a text-to-image generative model. We introduced two feedback mechanisms to find prompts informed by both the given model and the desirable target image distribution. Evaluations on a diverse set of tasks and distribution shifts show the effectiveness of the proposed closed-loop approach in comparison to open-loop ones. Below we discuss some of the limitations of our work:
  • Label shift: In this work, we focus on generating novel images. However, some distribution shifts can also change the label distribution, e.g., for depth estimation, changing from indoor to outdoor scenes would result in a shift in depth maps. One possible approach could be learning a generative model over the label space (Le Moing et al., 2021) to control the generation in both the label and image space.
  • Computational cost: Estimating the gradient of the loss in Eq. 3 requires backpropagation through the denoising process of the diffusion model, which can be computationally demanding. Using approaches that reduce the number of denoising steps (Song et al., 2020; Luo et al., 2023) may be able to reduce this computational cost.
@@ Appendix A Appendix Outline
  • Appendices B and C describe additional implementation details for our classification and depth estimation experiments, respectively.
  • Sec. B.6 describes an image guidance mechanism using Textual Inversion (Gal et al., 2022) and compares it with the CLIP guidance mechanism, on Waterbirds.
  • Sec. C.2 provides additional quantitative results. See Sec. C.2.1 for comparisons with other data augmentation baselines for depth estimation and Sec. C.2.2 for comparisons with domain generalization baselines on iWildCam. See Sec. C.2.3 for results on training with more synthetic data.
  • Secs. B.4.3, B.5.4 and C.3 provide qualitative generations from all the Adversarial Prompts and Guided Adversarial Prompts used in the Waterbirds, iWildCam, and depth estimation experiments.
    • Additionally, for depth estimation, we provide a qualitative comparison of Adversarial Prompts generations optimized on different models (UNet (Ronneberger et al., 2015), DPT (Ranftl et al., 2021)). For iWildCam, we also provide additional results using a ViT-B-16 (Dosovitskiy et al., 2020) instead of a ResNet50 (He et al., 2016b).
  • Sec. C.4 provides additional analysis on the depth estimation experiments:
@@
Appendix B Implementation Details for Classification Tasks
B.1 Training data generation
@@
As mentioned in main paper Sec. 3.1, for semantic classification tasks, we utilize the foreground object masks and use an in-painting technique proposed in Lugmayr et al. (2022) that preserves the masked region throughout the denoising process. In this section, we briefly describe this procedure and refer the reader to the original work for more details.
@@
However, because we are using Stable Diffusion Rombach et al. (2022), the denoising process is done in latent space (using an encoder $\mathcal{E}$), not pixel space. This means that to apply inpainting, we must resize the mask $m$ to the latent space dimensions, and apply the above-described procedure in the latent space: $\tilde{z}_{t}\leftarrow m_{z}\odot z^{\mathrm{orig}}_{t}+(1-m_{z})\odot\tilde{z}_{t}$, where $z^{\mathrm{orig}}_{0}=\mathcal{E}(x^{\mathrm{orig}})$ and $z^{\mathrm{orig}}_{t}$ is its corresponding noised version. While this procedure usually performs well in preserving the original region of interest, we also paste the original masked region in the pixel space to obtain the final sample $\tilde{x}=m\odot x^{\mathrm{orig}}+(1-m)\odot\tilde{x}_{0}$.
SDEdit Meng et al. (2021a).
In addition to inpainting, depending on the setting, we also use SDEdit Meng et al. (2021a), a mechanism available to all diffusion models that allows to use an initial image to condition the generation of new images to be closer to the initial image. The mechanism is parametrized by the SDEdit strength $s$, which indicates the extent by which the model can deviate from the original image.
@@
B.2 ALIA
Here, we give more details on the ALIA Dunlap et al. (2023) (Automated Language-guided Image Augmentation) baseline method, which aims at generating images targeting a particular test distribution similar to our guidance mechanism (main paper Sec. 3.3).
Given exemplar images from the test distribution, ALIA first captions each image using the BLIP Li et al. (2022) captioning model. Then, it uses the GPT-4 OpenAI (2023) LLM to summarize these captions into a list of domains asking it to produce descriptions that are agnostic to the class information. Dunlap et al. (2023) then use these prompts to generate additional training data. In order to preserve the original class information in their generations, they use SDEdit Meng et al. (2021a) or Instruct Pix2Pix Brooks et al. (2023). We refer the original paper for further implementation details. Below, we summarize resulting prompts we use for comparison in our results.
For Waterbirds Sagawa et al. (2019), we found that removing the prefix “a photo of a {class name}" from the original prompts when using the inpainting technique (Sec. B.1) to work slightly better for both the ALIA baseline and our CLIP text guidance (main paper Eq. 2). We, therefore, use the following prompts:
@@ B.2 ALIA
For iWildCam Beery et al. (2021), we keep the original prompts intact:
@@
B.3 General implementation details
Table 3: Generation and optimization parameters for Adversarial Promptsand Guided Adversarial Promptsfor classification tasks in Sec. 4.1.
Waterbirds 1. 5 7.0 5 1000/1000 20/0
iWildCam 0.8 5 5.0 10 2000/10000 0/10
@@ B.3 General implementation details
We report our generation parameters in Tab. 4. We use the DDIM Song et al. (2020) scheduler. We generate 384x384 resolution images. Those parameters were chosen based on visual inspection, ease of optimization and downstream performance (validation accuracy). Training data. After generation, ALIA’s method consists of an additional filtering step to remove “bad” generations. This step relies on using a pretrained model to measure confidence on the generated images. However, given our method creates images that are adversarial to an iWildCam pretrained model, the filtering part of ALIA’s pipeline is not usable on our data. Thus, to keep things comparable, we decided not to apply filtering both our method generated data and ALIA’s generated data. However, it must be noted that Dunlap et al. (2023) only reports a 2% absolute accuracy drop between fitlering and no filtering on iWildCam (1.4% on Waterbirds), thus we do not expect a big difference in performance with ALIA’s reported results and our results.
@@
We report our training parameters in Tab. 5. We use ALIA’s codebase to finetune our models, which ensures fair comparison to the ALIA baselines. For everything except the generated data, the settings are the same as in ALIA. For both datasets, the starting model is a ResNet50 He et al. (2016a) model, pretrained on ImageNet Deng et al. (2009).
@@
B.4 Waterbirds
B.4.1 Dataset details
Fig. 7 demonstrates the shift between train and test distributions in the Waterbirds dataset Sagawa et al. (2019). We follow the setting suggested in Dunlap et al. (2023) and use 1139 images as $\mathcal{D}_{\mathrm{tr}}$, where waterbirds appear only on water background and landbirds on land background. We add additional 839 examples either from the original dataset, where waterbirds appear only on land background and landbirds on water background ("Real OOD data"), or generated by Stable Diffusion with prompts obtained by one of the methods. For the data-efficiency plots (e.g., Fig. 3) we reduced the number of added examples by a factor of {1/2, 1/4, 1/8, 1/16}.
Since the original Waterbirds dataset does not provide masks for the exact generated images, we used the SAM Kirillov et al. (2023) segmentation model to obtain bird segmentation masks for training images. We use these masks to condition the generative model on the class by using inpainting as described in Sec. B.1.
B.4.2 Implementation Details
@@
For adversarial feedback, we use the model trained only using the original training data $\mathcal{D}_{\mathrm{train}}$ with complete spurious correlation. It is taken from ALIA checkpoints [^3] https://api.wandb.ai/files/clipinvariance/ALIA-Waterbirds/y6zc932x/checkpoint/ckpt-Waterbirds-none-filtered-resnet50-1-0.001-0.0001/best.pth. As the task is the binary classification, we use the cross-entropy loss for the opposite class as the adversarial loss: $\mathcal{L}_{\mathrm{adv}}(\tilde{x},y)=\mathcal{L}_{\mathrm{x-ent}}(f(\tilde{x}),1-y)$, assuming $y\in\{0,1\}$. This is equivalent to the negative cross-entropy loss referred in the text. We find four prompts per each class, i.e., eight prompts in total. Each prompt is composed of five new learnable tokens. We perform adversarial optimization for 1000 steps with learning rate 1e-3 using Adam Kingma & Ba (2017). We use five denoising steps during adversarial optimization and generate images for training with 15 steps. We do not use SDEdit for Waterbirds. See Tab. 4 for summary.
@@
For Waterbirds, we use CLIP text guidance by encoding each of ALIA’s summarized prompts (see Sec. B.2) with the CLIP text encoder as described in main paper Sec. 3.3. In addition, we renormalize the averaged target text embedding to have the norm equal to the mean norm of the original prompts, and use the resulting vector as the target $e_{\mathrm{t}}$. We use $l_{2}$ guidance loss: $\mathcal{L}_{\mathrm{t}}(E_{\mathrm{t}}(c_{w}),e_{\mathrm{t}})=\|E_{\mathrm{t}}(c_{w})-e_{\mathrm{t}}\|_{2}^{2}$. We use $\lambda_{\mathrm{t}}=20$ and $\lambda_{\mathrm{i}}=0$ (i.e., no image guidance).
B.4.3 Additional Qualitative Results.
@@
B.5 iWildCam
B.5.1 Dataset details.
The original iWildCam Beery et al. (2021) dataset is subsampled to create a 7-way classification task (background, cattle, elephant, impala, zebra, giraffe, dik-dik). The training set has 6,000 images with some classes having as few as 50 images per example. There are 2 test locations that are not in the training or validation set. Additionally, given $h$, the hour at which an image was taken, we define an image to be during “daytime” if $9\leq h\leq 17$, and “nighttime” if $h\leq 5\lor h\geq 20$. As said in main paper Sec. 4.1, for image CLIP guidance, we separate the target test locations into four groups (location={1,2}, time={daytime, nighttime}). We provide visualisation of the test locations (at day & night) in Fig. 10. For more details on the iWildCam subset construction, we refer to Dunlap et al. (2023) Section 8.3. For inpainting, the object masks are obtained from MegaDetector Beery et al. (2019).
@@
B.5.2 Alignment collapse solution for iWildCam.
@@ B.5.3 Implementation details
We use the “entropy” loss, described previously. For adversarial prompts, we train for 2000 steps. For guided adversarial prompts, we use CLIP guidance coefficient with $\lambda_{i}=10$ and and $\lambda_{\mathrm{t}}=0$ (i.e., no text guidance). We train for a total of 10000 gradient steps. However, we don’t optimize the adversarial loss for the first 2000 steps to allow the prompt to first converge to the target distribution region. One run takes about 3 hours, using one A100 GPU with 80GB of RAM, on our internal cluster.
@@
The generation settings are the same as the ones used during adversarial optimization. For each target domain-guided adversarial prompt, (i.e. location & time of the day), the source images (used to condition the generation with an object mask and through SDEdit Meng et al. (2021a)) are only images that match the time of the day of the target domain used during generation. Furthermore, for each prompt, we only generate one image per source image.
@@
Our iWildCam pretrained model is taken from ALIA checkpoints [^5] https://api.wandb.ai/files/clipinvariance/ALIA-iWildCamMini/brr7b3ks/checkpoint/ckpt-iWildCamMini-randaug-filtered-resnet50-0-0.001-0.0001/best.pth. ALIA trains the model from “scratch” (i.e. the model has never seen iWildCam data), for 100 epochs, on the combination of real + generated data. For our method, given we optimize the prompts based on a finetuned model feedback, it may not make as much sense to train the model from "scratch". Thus, we also introduce the variant where the iWildCam pretrained model is finetuned on the combination of real + generated data for 20 epochs, where finetuning means that every layer, except the last, is frozen.
@@
B.5.4 Additional Qualitative Results.
@@
B.5.5 Using ViT model.
In Fig. 14, we repeat the iWildCam experiment from the main paper (Fig. 4) with a ViT-B-16 Dosovitskiy et al. (2020) model. Additionally, we provide qualitative results for generations from adversarial prompts optimized with a ViT-B-16 model in Fig. 15.
@@
B.6 Image Guidance using Textual Inversion
In addition to the CLIP image guidance introduced in main paper Sec. 3.3, we also explore using Textual Inversion (TI) Gal et al. (2022) as an image guidance mechanism. Similar to the CLIP guidance, we use a few images $\{x_{j}\}$ from the target distribution. Now, instead of the similarity in the CLIP embedding space, we use the denoising loss between a generated image and one of the target images (see Eq. (2) in Gal et al. (2022)):